For a manufacturing industry, the number of industrial accidents averages three per week.

a.    Find the probability that at most five accidents will occur in a given week. 



b.    Find the probability that two accidents will occur in a given day.




1
Expert's answer
2021-09-07T19:21:11-0400

"x=number\:of\:accidents\:per\:week\:"


"Average\:number\:per\:week\left(\lambda \right)=3"


"The\:poisson\:distribution\:function=\frac{e^{-\lambda }\lambda \:^x}{x!}\:\:\:x=1,2,3...."


a. "P\left(x\le 5\right)=P\left(x=0\right)+P\left(x=1\right)+P\left(x=2\right)+P\left(x=3\right)+P\left(x=4\right)+P\left(x=5\right)"

"=0.8173"



b. The average number of accidents per day is given as:

"\:\lambda _d=\frac{3}{7}"

"y=number\:of\:accidents\:per\:given\:day"

"P\left(y=2\right)=\frac{e^{-\frac{3}{7}}\left(\frac{3}{7}\right)^2}{2!}=0.00001312"